1. (15 points) Your grandfather has offered you a choice of one of the theefollowing alternatives: $4500 now; $1000 at the beginning of each period for six years, or $8000 at the end of six years. Assuming you could earn 12% annually.
Required: (1) Calculate the Present Value of each choice (2) State which alternative should you choose?
Answer #1
1)
a) PV of $4,500 now = $        4,500
b) PV of $1,000 for BOY 1 to 6 (annuity due) = 1000*1.12*(1.12^6-1)/(0.12*1.12^6) = $        4,605
c) PV of $8,000 at EOY 6 = 8000/1.12^6 = $        4,053
2) Alternative [b] should be chosen as it has the highest
PV.